Theobromine is a bitter alkaloid of cacao plant. It is found in chocolate, leaves of tea plant and cola nut. How many grams of theobromine (C7H8N4O2) should be dissolved in 375g water to make 3.75 m theobromine solution?


1
Expert's answer
2021-03-08T06:10:37-0500

C7H8N4O2 = 180.164 "(\frac{g}{mole})"


(3.75 "\frac{mole}{kg}" )*(0.375 kg)*(180.164 "(\frac{g}{mole})" ) = 253 g of C7H8N4O2